# Conceptualization, Writing—original draft

### v2

@@ 4.4 Ablation study
Table 5: Ablation study of the effectiveness of the components of the NBED on BSDS dataset. All results are computed with a single scale test. Local-Feat: only use location feature, Seman-Feat: only use semantic feature, Full: Our proposed method (location feature + semantic feature), CA: Channel Attention [32], ASPP: Atrous Spatial Pyramid Pooling [11], CoFusion: Context-aware fusion block [13].
@@
In the first set of experiments, we test the effect of various encoders. VGG16 is a widely used encoder in edge detection. when VGG16 is used as the encoder, the ODS of the model is 0.797, which is similar to RCF [18] and LDC [30]. When only the location feature encoder of NBED is used, the ODS of the model is only 0.688, which is far worse than 0.823, the ODS semantic feature encoder. And these three encoders are far less effective than our NBED, which can achieve an ODS of 0.838. A more intuitive difference can be found from Fig. 5. The location feature encoder can preserve the location information of the edge, but due to the lack of semantic information guidance, the edge map is mixed with a lot of noise. On the contrary, the semantic feature encoder can well suppress the expression of noise by using semantic information, but the lack of accurate location information produces aliasing in the upsampling process, which makes the edges rough. The results of VGG16 encoder look a little better, which is due to the ability of VGG16 to extract both location and semantic information. However, due to the lack of feature extraction ability of VGG16, it is difficult to meet the needs of edge detection, and ODS is much lower than NBED. Another point worth noting is that although the edges generated by semantic feature encoder are not as good as VGG16 in visualizing, it is more accurate after NMS.
@@
In addition, we verify the effectiveness of the feature enhancement and fusion modules in previous methods [11, 32, 13] on our NBED. When the CA module and ASPP module are added, the accuracy of the model does not increase but decreases, which indicates that when the encoder is powerful enough, these additional feature enhancement modules not only cannot further enhance the features, but will have a negative impact. Using the feature fusion module CoFusion, the model accuracy is almost unchanged, which indicates that the shallowest features of NBED have obtained sufficient semantic information, and no additional gains can be obtained by fusing features at other scales.

### v3

@@ 4.4 Ablation study
Table 5: Ablation study of the effectiveness of the components of the NBED on BSDS dataset. All results are computed with a single scale test. Local-Feat: only use location feature, Seman-Feat: only use semantic feature, Full: Our proposed method (location feature + semantic feature), CA: Channel Attention [su2021pixel], ASPP: Atrous Spatial Pyramid Pooling [he2022bdcn], CoFusion: Context-aware fusion block [huan2021unmixing].
@@
In the first set of experiments, we test the effect of various encoders. VGG16 is a widely used encoder in edge detection. when VGG16 is used as the encoder, the ODS of the model is 0.797, which is similar to RCF [liu2019richer] and LDC [soria2022ldc]. When only the location feature encoder of NBED is used, the ODS of the model is only 0.688, which is far worse than 0.823, the ODS semantic feature encoder. And these three encoders are far less effective than our NBED, which can achieve an ODS of 0.838. A more intuitive difference can be found from Fig. 5. The location feature encoder can preserve the location information of the edge, but due to the lack of semantic information guidance, the edge map is mixed with a lot of noise. On the contrary, the semantic feature encoder can well suppress the expression of noise by using semantic information, but the lack of accurate location information produces aliasing in the upsampling process, which makes the edges rough. The results of VGG16 encoder look a little better, which is due to the ability of VGG16 to extract both location and semantic information. However, due to the lack of feature extraction ability of VGG16, it is difficult to meet the needs of edge detection, and ODS is much lower than NBED. Another point worth noting is that although the edges generated by semantic feature encoder are not as good as VGG16 in visualizing, it is more accurate after NMS.
@@
In addition, we verify the effectiveness of the feature enhancement and fusion modules in previous methods [he2022bdcn, su2021pixel, huan2021unmixing] on our NBED. When the CA module and ASPP module are added, the accuracy of the model does not increase but decreases, which indicates that when the encoder is powerful enough, these additional feature enhancement modules not only cannot further enhance the features, but will have a negative impact. Using the feature fusion module CoFusion, the model accuracy is almost unchanged, which indicates that the shallowest features of NBED have obtained sufficient semantic information, and no additional gains can be obtained by fusing features at other scales.